consider two state system in thermal equilibrium, where the upper state has an energy 3.09 kj/mol higher than the lower state. if we have total population of thousand molecules distributed between these states, at what temperature (in c) will 15.6% of the molecules be in the upper state?


1
Expert's answer
2021-02-19T10:31:23-0500

Answer

Population is given by

"\frac{N_{upper}}{N_{lower}}=e^{-(\frac{E_1-E_2}{kT})}"

Putting all values as above theory of question

"\frac{15.6}{100\times1000}=e^{-(\frac{3090\times10^3}{8.314\times T})}"

"\frac{10^6}{156}=e^{(\frac{3090\times10^3}{8.314\times T})}"

Taking natural logrithm

And temperature

"T=\frac{3090\times1000}{8.314\times8.8}=42234K"